(1 point) Consider the linear system 3=[} }); a. Find the eigenvalues and eigenvectors for the coefficient matrix. EL and 12I AM REALLY STRUGGLING ON THIS PROBLEM PLEASE HELP ME CORRECT AND NEAT WORK IS MUCH APPRECIATED THANKS
Answer #1

a.

\begin{vmatrix} 3-x &2 \\ -5&-3-x \end{vmatrix}=-(3-x)(3+x)+10=x^2+1\\ So,eigen \ values\ are \ given\ by\ roots\ of\ x^2+1=0\\ i.e. \ i ,-i\\ now,for\ x_1=-i : \begin{bmatrix} 3+i &2 \\ -5&-3+i \end{bmatrix}\sim \begin{bmatrix} 1 & \frac{1}{5}(3-i)\\ 0&0 \end{bmatrix} So,take \ eigen \ vector \ v_1= \begin{pmatrix} \frac{1}{5}(-3+i)\\ 1 \end{pmatrix}t\\ then,for\ x_2 =i :\begin{bmatrix} 3-i &2 \\ -5&-3-i \end{bmatrix}\sim \begin{bmatrix} 1 & \frac{1}{5}(3+i)\\ 0&0 \end{bmatrix} So, take\ the\ eigen\ vector\ v_2=\begin{pmatrix} -\frac{1}{5}(3+i)\\ 1 \end{pmatrix}t

b.the \ general\ solution\ is,\\ y(t)=c_1v_1e^{x_1t}+c_2v_2e^{x_2t}=c_1\begin{pmatrix} \frac{1}{5}(-3+i)\\ 1\\ \end{pmatrix}e^{-it}+c_2\begin{pmatrix} -\frac{1}{5}(3+i)\\ 1\\ \end{pmatrix}e^{it}\\ Now,y(t)=\begin{pmatrix} y_1(t)\\ y_2(t) \end{pmatrix}\\ given, that \ y(0)=\begin{pmatrix} y_1(0)\\ y_2(0) \end{pmatrix}=\begin{pmatrix} 5\\ -5 \end{pmatrix}\\ \therefore ,\begin{bmatrix} \frac{1}{5}(-3+i)&-\frac{1}{5}(3+i) \\ 1&1 \end{bmatrix}\begin{pmatrix} c_1\\ c_2 \end{pmatrix}=\begin{pmatrix} 5\\ -5 \end{pmatrix}\\\Rightarrow \begin{pmatrix} c_1\\ c_2 \end{pmatrix}=\begin{pmatrix} -\frac{5i}{2} &\frac{1}{2}-\frac{3i}{2} \\ \frac{5i}{2}& \frac{1}{2}(1+3i) \end{pmatrix}\begin{pmatrix} 5\\ -5 \end{pmatrix}=\begin{pmatrix} -\frac{5}{2}-5i\\ -\frac{5}{2}+5i\\ \end{pmatrix}\\ \therefore y_1(t)=(-\frac{5}{2}-5i)(\frac{1}{5}(-3+i))e^{-it}+(-\frac{5}{2}+5i)(-\frac{1}{5}(3+i))e^{it} \\=\frac{5}{2}(1+i)e^{-it}+\frac{5}{2}(1-i)e^{it}=5cos(t)+5sin(t) \\y_2(t)=(-\frac{5}{2}-5i)e^{-it}+(-\frac{5}{2}+5i)e^{it}=-5cos(t)-10sin(t)